# Do Mice Grok? Glimpses of Hidden Progress During Overtraining in Sensory Cortex

### v1

@@ 1 Introduction
In deep learning, this has changed with the discovery of the grokking phenomenon (Power et al., 2022), whereby neural networks first fit their training data to high accuracy, then many epochs of training later (“overtraining”), abruptly generalize. As a phenomenon that epitomizes how neural networks can behave in sharp and unexpected ways, it has been under intense study (Nanda et al., 2023; Liu et al., 2022a; Varma et al., 2023). Emerging theories attribute grokking to causes as diverse as weight decay (Nanda et al., 2023), initialization weight norm (Liu et al., 2022a), inductive bias, (Varma et al., 2023), and more. Much recent theoretical work attempts to unify these as positing that generalization occurs due to delayed feature learning taking place when the network has already achieved ceiling accuracy on the training task (Lyu et al., 2023; Kumar et al., 2023; Mohamadi et al., 2024).
@@
In psychology, there exists a long line of work on learning tasks at all levels of expertise, including after task mastery (Ericsson et al., 1993; Newell and Rosenbloom, 1981; Fitts and Posner, 1967; Mackintosh, 1969; Richman et al., 1972; Mead, 1973). Many of these works attempt to address the intuitive observation that human experts in a variety of domains continue to learn on tasks at which they have achieved near-ceiling performance.
@@
An understanding of learning dynamics during overtraining, however, is broadly missing at the level of experimental neuroscience. This is largely because there are several difficulties in testing the hypothesis that representations continue to evolve during overtraining in cortex. First, for many animals, recording neurons is invasive and risks damaging the animals, so recordings are typically only taken at the end of training to study the end-time learned representation (Yuste, 2015; Kim et al., 2016). Second, experiments in systems neuroscience do not typically design stimuli as “training” and “test” in the way datasets are constructed in deep learning, so making claims about learning and generalization is not possible. We elaborate on further difficulties, as well as the connection of our phenomenon of interest to the related phenomenon of representational drift, in Appendix A.5.
@@
Here, we revisit the neural data from a recent work with precisely the desirable qualities above: Berners-Lee et al. (2023), who study neural activity during learned odor discrimination task in mouse piriform cortex. The mice are trained to discriminate one target odor from hundreds of nontarget odors, then continue to be trained for weeks on the same task after reaching behavioral mastery (“overtraining”), with neural firing rate data recorded during this time. After overtraining, they are given held-out test examples. Our starting point in this work are two empirical observations in (Berners-Lee et al., 2023) that we seek to model and explain.
@@
  • Decoding accuracy of training labels from population activity increases throughout overtraining, despite no visible changes in behavior.
@@
  • We find that the margin of the maximum margin classifier at each day of overtraining is increasing in mouse cortex, implying that points (odor trials) sufficiently close to the optimal decision boundary may be classified incorrectly at the beginning of overtraining, but correctly at the end, despite no change in behavior during this time.
  • We construct a biologically faithful synthetic model of piriform cortex performing this task exhibits these properties, as well as the grokking phenomenology, out-of-the-box. We can interpret the resulting dynamics and trace the abrupt learning to margin maximization.
@@
We also note that (Berners-Lee et al., 2023) are not the first to observe such effects experimentally: similar effects in sharpening of already-learned odor representations was also observed by Shakhawat et al. (2014); Kadohisa and Wilson (2006), so that the phenomena we seek to model in this work are likely to be robust at least in the setting of mouse olfaction. We note our neural data reanalysis is observational, and outline associated limitations in Section 5. We also include detailed experimental proposals in which our hypotheses can be tested directly in Appendix Section A.2.
@@ 2 Related Work
Lazy/rich learning and margin maximization. We cast our ideas in the language of “lazy” and “rich” regimes of representation learning, introduced in machine learning (Chizat et al., 2019; Jacot et al., 2018). Such characterization of learning regimes has become increasingly popular in computational neuroscience (Chizat et al., 2019; Farrell et al., 2023; Flesch et al., 2021, 2022; Ito and Murray, 2023). In machine learning, lazy learning refers to a trained neural network being well approximated by a kernel method in its initial neural tangent kernel (NTK) and the rich regime is when this approximation breaks down as the network learns task-relevant features and its weights move far from initialization. In classification problems, margin maximization can be one type of rich learning Mohamadi et al. (2024); Matyasko and Chau (2017), where the margin of a classifier is defined as the distance from the decision boundary to the nearest data point. There are some subtleties around defining margin maximization in a setting where a decoder is being retrained on an evolving feature map; we state a precise working definition in Section 3.1.2.
Grokking. Much theoretical work has been done to understand the grokking phenomenon, discovered by (Power et al., 2022), in which task-specific representation learning occurs long after training accuracy has saturated (Nanda et al., 2023; Barak et al., 2022; Liu et al., 2022b). Current theories explaining grokking are varied: (Liu et al., 2022a) claim it is due to weight norm at initialization, (Nanda et al., 2023) suggest it is driven by weight decay, (Varma et al., 2023) suggest it is due to competition between memorizing and generalizing neural circuits. A flurry of recent theoretical papers unify these views this as late-time feature learning, where networks begin lazy and then abruptly transition into the rich regime much later during training (Kumar et al., 2023; Lyu et al., 2023). Recent empirical evidence has accumulated in support of this perspective (Edelman et al., 2024; Clauw et al., 2024; Mohamadi et al., 2024), where rich learning often takes the form of margin maximization in classification tasks. For instance, Morwani et al. (2023) and Mohamadi et al. (2024) show that on certain classes of tasks, margin maximization during overtraining provably causes grokking.
@@ 3 Revisiting Neural Data From Overtraining
Figure 1: (a) Behavior of mice on a binary discrimination task of odors, where mice indicate their selected choice by licking left or right. Y-axis is fraction of licks left vs right on a given day. Day 8 is the first day of overtraining. Correct lick for non-target is left, and right for target. The mice can discriminate near-perfectly as overtraining begins. Reproduced with modification from Berners-Lee et al. (2023). (b) Increase in decoding accuracy from 10-fold linear discriminant analysis for each session. Shaded colored lines in background show standard errors for each mouse.
@@ 3.1 Setup
Berners-Lee et al. (2023) collected tetrode recordings from posterior piriform cortex in adult mice during overtraining on a binary odor discrimination task (Figure 1). The posterior piriform cortex (PPC) is a key brain region involved in olfactory processing (Blazing and Franks, 2020; Srinivasan and Stevens, 2017). An odor is an $n$-hot vector of length $k$, with $n=3,k=13$. The “target” odor is uniquely defined as consisting of three specific odorants, where non-targets never contain any of these three. The discrimination task mice are trained on is to distinguish the target from the nontarget odors. After an initial training period of 8 days, the mice master the task, reaching ceiling behavioral performance (see Figure 1a), yet training continues in the same way for a further 18 days (“overtraining”). This period is when firing rate data is recorded. After this overtraining period is complete, the mice are exposed to held-out odors (“test/probe” examples), which share one odorant with the target odor and therefore are more difficult to correctly classify.
@@ 3.1.1 Representations of Target/Nontarget Odors Separate During Overtraining
The measurement of such representational similarity metrics is standard in neuroscience (Kriegeskorte et al., 2008), and similar to common interpretability techniques in deep learning that involve probing for directions in activation space (Bau et al., 2019; Olah et al., 2020; Elhage et al., 2021; Zhang and Nanda, 2023) and comparing them, for instance by computing their dot product. The continued separation of class representations between the target and nontarget odors in Figure 2 provides evidence that, despite no visible changes in behavior, there is continued representation learning in mouse PPC during overtraining on this task. Such continued learning may aid in generalization on unseen odors, as indeed (Berners-Lee et al., 2023) find empirically to be the case.
Figure 2: Representational similarity matrices plotting average pairwise correlation between target and nontarget population responses from piriform cortex on the first (top) and last (bottom) days of overtraining. We drop the last few days of data for Mouse T due to inconsistencies in data; methodological details in A.1. We compare the separations (anticorrelation) above to that to random/ablated baselines in Appendix A.4, finding them significantly larger.
@@
Figure 3: Neural representations (projected onto the top 2 principal components) of target and nontarget odors on the first (top) and last (bottom) day of overtraining for each mouse, showing qualitative separation during overtraining measured quantitatively in the representational similarity matrices in Figure 2. Each column corresponds to one mouse.
@@ 3.1.2 Representational Changes Lead to an Improved Maximum Margin Classifier
We begin by clarifying two distinct notions of margin maximization that exist when a decoder is trained on an evolving hidden representation. Consider a feature map $\phi_{t}(x)$ defined by a population of neurons at time $t$, as well as a decoder $d_{n}$ trained for $n$ steps on this fixed feature map to predict class labels from firing activity. First, the margin of $d_{n}(\phi_{t})$ will increase in $n$, even as $t$ remains fixed and the features remain unchanged (Soudry et al., 2017). Second: if, instead, $n$ remains fixed, then the only way for $d_{n}(\phi_{t}(x))$ to increase is due to a task-aligned improvement in the features—rich learning in $\phi_{t}(x)$. This second case where margin maximization happens because hidden features improve in task-relevant directions will be the notion of margin maximization we will use.
Formally, let $\phi_{t}(x)\in\mathbb{R}^{N}$ be the representation at time $t$ for data $x$, and let $w(t)$ be the parameters of the max margin (support vector machine, SVM) solution to the label classification problem with $\phi_{t}(x_{i})$ as input point $i$ and $y_{i}$ as label, then the margin $\mathcal{M}(t)$ takes the form:
@@
Note that this margin is a function of time, since the neural representation $\phi_{t}(x)$ changes over time, and intuitively represents “the margin of the max margin classifier on a given set of hidden features.” In this sense, our hypothesis is that the evolution on odor representations takes the form of (max) margin maximization.
@@
Therefore, PPC representations evolve to increase the margin of the maximum margin classifier trained on those representations, similar to how several researchers (Morwani et al., 2023; Mohamadi et al., 2024; Lyu and Li, 2019) find that deep neural networks which “grok” during overtraining on classification tasks are implicitly driven to do so by maximizing margin.
This margin-maximization perspective is provocative not only because it provides a geometric and intuitive picture of continued learning in this setting, but because it is suggestive of why certain generalizing solutions are learned. For instance, (Morwani et al., 2023; Mohamadi et al., 2024; Lyu and Li, 2019) prove in various settings that the reason that neural networks often solve particular tasks (eg. modular arithmetic) using very specific types of features (eg. Fourier features) (Nanda et al., 2023), is because the regularized optimization trajectories are implicitly biased towards margin-maximizing solutions. We emphasize that since test data (unseen test odors) is missing during the training period, we cannot make claims about whether the cortical representation begins lazy or rich (ie. evolve significantly during the initial 7-day learning period compared to initialization), and this is an important avenue for future work and limitation of our reanalysis. While the optimization algorithms at play in brains are poorly understood, it is speculatively possible that some form of “implicit-bias” towards generalizing solutions may also be the reason that overtrained mice do not just memorize their training odors, but learn representations that are helpful on difficult, held-out examples.
@@ 4.1 A synthetic model of mouse piriform cortex
Setup. We construct a synthetic version of the odor discrimination task under biological constraints. Piriform cortex is known to receive input from mouse glomerular channels in a relatively random and disperse manner (Blazing and Franks, 2020). We take our earlier definition of odors as $n$-hot vectors of odorants (chemicals) in $\{0,1\}^{k}$ that we pass through a random projection, inspired by the functional role of the glomeruli (Blazing and Franks, 2020). Then, we train a one-hidden layer multilayer perceptron (MLP) on these embeddings to classify the one target odor from 200 non-target odors. We use a ReLU nonlinearity to simulate non-negative firing rates. We also hold out 20 test examples that share an odorant (a single entry in the bit string) with the target. These examples are are thus closer in both Hamming and projected space to the target, and therefore harder to discriminate. We train the MLP with vanilla gradient descent on a cross-entropy loss, as is standard in classification.
@@
Mouse behavior and neural dynamics are captured by an overtrained MLP. This setup replicates the behavior and neural correlates of the mouse (Figure 5). The test loss (on the probe trials) continues to decrease for thousands of epochs after training loss plateaus at a low value, in line with a continued increase in the classifier margin. In this setting, the readout layer of our MLP acts as the “decoder” on the hidden layer representation, which are evolving to allow for an increased max-margin solution to be found. We also see a separation between the target and the nontarget representations emerge during training (up to epoch 1000). However, at epoch 1000, when training loss has converged, the target and probes are not yet easily separable. Strikingly, overtraining on the target-nontarget separation task improves performance on the probe odors, which are out-of-distribution from the training set. This happens because the network maximizes margin between the target and nontarget representations during the overtraining period after epoch 1000. This manifests as the red dot (target) gradually separating from the green cluster (test probes) during overtraining. There is an abrupt grokking-like transition in test accuracy (the fraction of probe trials correctly classified) during the overtraining period, Epochs 1000-9000, exactly as we see in Figure 6.
@@
Possible biological drivers of late time learning. Certain types of loss functions have been proven in a deep learning setting to drive late-time margin maximization (Soudry et al., 2017; Lyu and Li, 2019). Here, we speculate on biologically plausible components of the objective in cortex that may plausibly serve a similar role. A simple but plausible driver of late-time learning is the set of energetic/metabolic constraints on biological organisms. Such constraints been suggested to give rise to a variety of neural phenomena (Sterling and Laughlin, 2015), for example disentangled representations (Pehlevan et al., 2017; Olshausen and Field, 1996; Plumbley, 2003; Whittington et al., 2023). Pertinently, such constraints, in the form of weight decay in machine learning, are often the driving force for grokking in some settings (Nanda et al., 2023; Liu et al., 2022a). These constraints force neural networks to deviate from using a lazy (kernel) method, and force the weights to move in task-relevant directions (Lyu and Li, 2019; Kumar et al., 2023). Another possibility in (Berners-Lee et al., 2023) is a very small supervised error signal from occasional mistakes made by the mice for idiosyncratic reasons that could plausibly be driving late-time learning (See Figure 1).

### v2

@@ 1 Introduction
In deep learning, this has changed with the discovery of the grokking phenomenon (Power et al., 2022), whereby neural networks first fit their training data to high accuracy, then many epochs of training later (“overtraining”), abruptly generalize. As a phenomenon that epitomizes how neural networks can behave in sharp and unexpected ways, it has been under intense study (Nanda et al., 2023; Liu et al., 2022a; Varma et al., 2023). Emerging theories posit that generalization occurs due to delayed feature learning taking place when the network has already achieved ceiling accuracy on the training task (Kumar et al., 2023; Lyu et al., 2023).
@@
In psychology, there exists a long line of work on learning tasks at all levels of expertise, including after task mastery (Ericsson et al., 1993; Newell and Rosenbloom, 1981; Fitts and Posner, 1967; Mackintosh, 1969; Richman et al., 1972; Mead, 1973). Many of these works attempt to address the intuitive observation that human experts in a variety of domains continue to learn on tasks at which they have achieved near-ceiling performance. There has also been much work on using deep learning to model perceptual learning tasks, to study the degree to which biological and artificial neural networks exhibit similar behavior on this class of tasks (Wenliang and Seitz, 2018; Bakhtiari, 2019; Yashar and Denison, 2017).
@@
The nature of representational changes at play during overtraining on learned tasks have been tackled at a fine grained (theorems proven about certain classes of deep networks) and coarse grained (qualitative explanations in psychology) levels. Such an understanding, however, is broadly missing at an intermediate level: that of experimental neuroscience. This is because there are several difficulties in naively testing the hypothesis that representations continue to evolve during overtraining in cortex. First, for many animals, recording neurons is invasive and risks damaging the animals, so recordings are typically only taken at the end of training to study the end-time learned representation (Yuste, 2015; Kim et al., 2016). Second, experiments in systems neuroscience do not typically design stimuli as “training” and “test” in the way datasets are constructed in deep learning, so making claims about learning and generalization is not possible. We elaborate on further difficulties, as well as the connection of our phenomenon of interest to the related phenomenon of representational drift, in Appendix A.3.
@@
Here, we revisit the neural data from a recent work with precisely the desirable qualities above: (Berners-Lee et al., 2023), who study neural activity during learned odor discrimination task in mouse piriform cortex. The mice are trained to discriminate one target odor from hundreds of nontarget odors, then continue to be trained for weeks on the same task after reaching behavioral mastery (“overtraining”), with neural firing rate data recorded during this time. After overtraining, they are given held-out test examples. Our starting point in this work are two empirical observations in (Berners-Lee et al., 2023) that we seek to model and explain.
@@
  • Decoding accuracy of training labels from population activity increases throughout overtraining.
@@
  • We find that the margin of the maximum margin classifier at each day of overtraining is increasing in mouse cortex, implying that points (odor trials) sufficiently close to the optimal decision boundary may be classified incorrectly at the beginning of overtraining, but correctly at the end, despite no outward change in behavior.
  • We construct a synthetic model of piriform cortex performing this task exhibits these properties, as well as the grokking phenomenology. We interpret the resulting dynamics and use targeted ablations to trace the continued learning to margin maximization.
@@
We also note that Berners-Lee et al. (2023) are not the first to observe such effects experimentally: similar effects in sharpening of already-learned odor representations was also observed by Shakhawat et al. (2014); Kadohisa and Wilson (2006), so that the phenomena we seek to model in this work are likely to be robust at least in the setting of mouse olfaction.
@@ 2 Related Work
Lazy/rich learning and margin maximization. We cast our ideas in the language of “lazy” and “rich” regimes of representation learning, introduced in machine learning (Chizat et al., 2019; Jacot et al., 2018). Such characterization of learning regimes has become increasingly popular in computational neuroscience (Chizat et al., 2019; Farrell et al., 2023; Flesch et al., 2021, 2022; Ito and Murray, 2023). In machine learning, lazy learning refers to a trained neural network being well approximated by a kernel method in its initial neural tangent kernel (NTK) and the rich regime is when this approximation breaks down as the network learns task-relevant features and its weights move far from initialization. In classification problems, margin maximization can be one type of rich learning (Mohamadi et al., 2024; Matyasko and Chau, 2017), where the margin of a classifier is defined as the distance from the decision boundary to the nearest data point. There are some subtleties around defining margin maximization in a setting where a decoder is being retrained on an evolving feature map; we state a precise working definition in Section 3.1.2.
Grokking. Much theoretical work has been done to understand the grokking phenomenon, discovered by Power et al. (2022), in which task-specific representation learning occurs long after training accuracy has saturated (Nanda et al., 2023; Barak et al., 2022; Liu et al., 2022b). Current theories explaining grokking are varied: Liu et al. (2022a) claim it is due to weight norm at initialization, Nanda et al. (2023) suggest it is driven by weight decay, Varma et al. (2023) suggest it is due to competition between memorizing and generalizing neural circuits. A flurry of recent theoretical papers unify these views this as late-time feature learning, where networks begin lazy and then abruptly transition into the rich regime much later during training (Kumar et al., 2023; Lyu et al., 2023). Recent empirical evidence has accumulated in support of this perspective (Edelman et al., 2024; Clauw et al., 2024; Mohamadi et al., 2024), where rich learning often takes the form of margin maximization in classification tasks. For instance, Morwani et al. (2023) and Mohamadi et al. (2024) show that on certain classes of tasks, margin maximization during overtraining provably causes grokking.
@@ 3 Revisiting Neural Data From Overtraining
Figure 1: (Left) Behavior of mice on a binary discrimination task of odors, where mice indicate their selected choice by licking left or right. Y-axis is fraction of licks left vs right on a given day. Day 8 is the first day of overtraining. Correct lick for non-target is left, and right for target. The mice can discriminate near-perfectly as overtraining begins. Reproduced with modification from (Berners-Lee et al., 2023). (right) Increase in decoding accuracy from 10-fold linear discriminant analysis for each session. Shaded colored lines in background show standard errors for each mouse.
@@ 3.1 Setup
Berners-Lee et al. (2023) collected tetrode recordings from posterior piriform cortex in adult mice during overtraining on a binary odor discrimination task (Figure 1). The posterior piriform cortex (PPC) is a key brain region involved in olfactory processing (Blazing and Franks, 2020; Srinivasan and Stevens, 2017). An odor is defined as an $n$-hot vector of length $k$, with $n=3,k=13$. In other words, a unique combination of three chemicals out of thirteen possible chemicals. The “target” odor is uniquely defined as consisting of three specific chemicals, where non-targets never contain any of these three. The discrimination task mice are trained on is to distinguish the target from the nontarget odors. After an initial training period of 8 days, the mice master the task, reaching ceiling behavioral performance (see Figure 1a), yet training continues in the same way for a further 18 days (“overtraining”). This period is when firing rate data is recorded. After this overtraining period is complete, the mice are exposed to held-out odors (“test/probe” examples), which share one odorant with the target odor and therefore are more difficult to correctly classify.
@@ 3.1.1 Representations of Target/Nontarget Odors Separate During Overtraining
The measurement of such representational similarity metrics is standard in neuroscience (Kriegeskorte et al., 2008), and similar to common interpretability techniques in deep learning that involve probing for directions in activation space (Bau et al., 2019; Olah et al., 2020; Elhage et al., 2021; Zhang and Nanda, 2023) and comparing them, for instance by computing their dot product. The continued separation of class representations between the target and nontarget odors in Figure 2 provides evidence that, despite no visible changes in behavior, there is continued representation learning in mouse PPC during overtraining on this task. Such continued learning may aid in generalization on unseen odors, as indeed Berners-Lee et al. (2023) find empirically to be the case.
Figure 2: Representational similarity matrices plotting average pairwise correlation between target and nontarget population responses from piriform cortex on the first (top) and last (bottom) days of overtraining. We drop the last few days of data for Mouse T due to inconsistencies in data; methodological details in A.1. We compare the separations (anticorrelation) above to that of random/ablated baselines in Appendix A.5, finding them significantly larger.
@@
Figure 3: Neural representations (projected onto the top 2 principal components) of target and nontarget odors on the first (top) and last (bottom) day of overtraining for each mouse, showing qualitative separation during overtraining measured quantitatively in the representational similarity matrices in Figure 2. Cluster centers plotted for ease of visual comparison.
@@ 3.1.2 Representational Changes Lead to an Improved Maximum Margin Classifier
We begin by clarifying a few distinct notions of margin maximization that apply when a decoder is trained on an evolving hidden representation.
  1. 1.
    A linear decoder is trained on a fixed feature representation $\phi$ for a fixed number of steps, which may or may not involve convergence of the decoder.
  2. 2.
    A linear decoder is trained on a fixed feature representation $\phi$ to convergence.
  3. 3.
    The feature representation $\phi_{t}$ changes over time, and at each time step $t$, a linear decoder is trained on the feature representation to convergence.
While the margin of the classifier defined by the decoder can increase as the decoder is trained for more steps on a fixed feature representation as in (1), if the margin of the max-margin (converged) classifier is increasing in (3), it implies the features are evolving to allow for a max-margin classifier with larger margin. This latter notion of margin-maximization that involves feature learning is the notion of margin maximization we will use.
We can define this mathematically as follows. Let $\phi_{t}(x)\in\mathbb{R}^{N}$ be the representation at time $t$ for data $x$, and let $w(t)$ be the parameters of the max margin SVM solution to the label classification problem with $\phi_{t}(x_{i})$ as input point $i$ and $y_{i}$ as label. Then the margin $\mathcal{M}(t)$ takes the form:
@@
In these terms, our hypothesis is that the evolution on odor representations takes the form of (max) margin maximization.
@@
Therefore, PPC representations evolve to increase the margin of the maximum margin classifier trained on those representations, similar to how many works find that deep neural networks which “grok” during overtraining on classification tasks are implicitly driven to do so by maximizing margin (Morwani et al., 2023; Mohamadi et al., 2024; Lyu and Li, 2019). This margin-maximization perspective is provocative not only because it provides a geometric and intuitive picture of continued learning in this setting, but because it is suggestive of why certain generalizing solutions are learned. For instance, Morwani et al. (2023); Mohamadi et al. (2024); Lyu and Li (2019) prove in various settings that the reason that neural networks often solve particular tasks (eg. modular arithmetic) using very specific types of features (eg. Fourier features in Nanda et al. (2023)), is because the regularized optimization trajectories are implicitly biased towards margin-maximizing solutions.
We emphasize that since test data (unseen test odors) is missing during the training period, we cannot make claims about whether the cortical representation begins lazy or rich (ie. evolve significantly during the initial 7-day learning period compared to initialization), and this is an important avenue for future work and limitation of our reanalysis. While the optimization algorithms at play in brains are poorly understood, it is speculatively possible that some form of “implicit-bias” towards generalizing solutions may also be the reason that overtrained mice do not just memorize their training odors, but learn representations that are helpful on difficult, held-out examples.
@@ 4.1 A synthetic model of mouse piriform cortex
Setup. We construct a synthetic version of the odor discrimination task that captures key computational considerations at an abstract level. We consider a more biologically plausible counterpart in Appendix A.7, finding the same phenomenology persists in the same way. We take our earlier definition of odors as $n$-hot vectors of odorants (chemicals) in $\{0,1\}^{k}$ that we pass through a random projection, inspired by the functional role of the glomeruli (Blazing and Franks, 2020), with $n=10,k=100$. We train a one-hidden layer multilayer perceptron (MLP) on these embeddings to classify the one target odor from 200 non-target odors. We use a ReLU nonlinearity to simulate non-negative firing rates. We also hold out 20 test examples that share an odorant (a single entry in the bit string) with the target. These examples are thus closer in both Hamming and projected space to the target, and therefore harder to discriminate. We train the MLP with vanilla gradient descent on a cross-entropy loss, as is standard in classification.
@@
Mouse behavior and neural dynamics are captured by an overtrained MLP. This simple model captures the key computational phenomena at play in piriform data that was the subject of analysis in Section 3. In Figure 5, we can see that the test loss (on the probe trials) continues to decrease after training loss plateaus at a low value, in line with a continued increase in the classifier margin. The readout layer of our MLP acts as the “decoder” on the hidden layer representation, which are evolving to allow for an increased max-margin solution to be found. We also see a separation between the target and the nontarget representations emerge during training (up to epoch 1000). However, at epoch 1000, when training loss has converged, the target and probes are not yet easily separable. Strikingly, overtraining on the target-nontarget separation task improves performance on the probe odors, which are out-of-distribution from the training set. This happens because the network maximizes margin between the target and nontarget representations during the overtraining period after epoch 1000. This manifests as the red dot (target) gradually separating from the green cluster (test probes) during overtraining. There is an abrupt grokking-like transition in test accuracy (the fraction of probe trials correctly classified) during the overtraining period, Epochs 1000-9000, exactly as we see in Figure 6.
We can test whether margin maximization is the causal factor in this simple model by ablating it and asking whether late-time test loss improvements vanish concomitantly. We do this by replacing the Cross-Entropy objective, which is known to drive late-time margin maximization (Lyu and Li, 2019; Lyu et al., 2023), with a Hinge loss which has a hard-margin objective instead. This means after the model achieves a margin $C$ on each training point, loss on that point vanishes, so the late-time loss incentive for margin maximization is removed. We find in this ablated setting margin saturates approximately when training loss does, and that test-loss saturates at this point in training as well, so that the grokking-like effect is ablated. Mathematical details and plots are in Appendix A.6, supporting the notion of a causal link.
@@
Possible biological drivers of late time learning. Certain types of loss functions have been proven in a deep learning setting to drive late-time margin maximization (Soudry et al., 2017; Lyu and Li, 2019). Here, we speculate on biologically plausible components of the objective in cortex that may plausibly serve a similar role. A simple but plausible driver of late-time learning is the set of energetic/metabolic constraints on biological organisms. Such constraints been suggested to give rise to a variety of neural phenomena (Sterling and Laughlin, 2015), for example disentangled representations (Pehlevan et al., 2017; Olshausen and Field, 1996; Plumbley, 2003; Whittington et al., 2023). Pertinently, such constraints, in the form of weight decay in machine learning, are often the driving force for grokking in some settings (Nanda et al., 2023; Liu et al., 2022a). These constraints force neural networks to deviate from using a lazy (kernel) method, and force the weights to move in task-relevant directions (Lyu and Li, 2019; Kumar et al., 2023). Another possibility in Berners-Lee et al. (2023) is a very small supervised error signal from occasional mistakes made by the mice for idiosyncratic reasons that could plausibly be driving late-time learning (See Figure 1).